given market demand Qd= 50 - P and market supply P= Qs + 5
A. find the market equilibrium price and quantity?
B. what would be the state of the market if market price was fixed at $ 25 per unit?
C. calculate and interpret price elasticity of demand at the equilibrium point
1
Expert's answer
2020-02-18T08:53:06-0500

A. Qd= 50 - P, P = Qs + 5 -> Qs = P - 5. 

The market equilibrium price and quantity are:

Qd = Qs,

50 - P = P - 5,

2P = 55,

P = $27.5.

Q = 27.5 - 5 = 22.5 units.

B. If market price was fixed at $25 per unit, then there will be a shortage (Qd > Qs).

C. Price elasticity of demand at the equilibrium point is:

"Ed = -1×\frac{27.5}{22.5} = -1.22,"

so the demand is elastic.